Please help me with solving this equation. I came up with the answer 5/6 but I am not sure that this is right. 6x + 4(3x – 2) = 6x + 2. Thank-you
Answer by stanbon(75887)

6x + 4(3x – 2) = 6x + 2
6x + 12x-8 =6x+2
12x = 10
x = 5/6
===========
Your answer looks good.
Cheers,
Stan H.